# Improving End-to-End Training of Retrieval-Augmented Generation Models via Joint Stochastic Approximation

### v1

@@ 5 Conclusion and Future Work
Remarkably, the potential advantage of the JSA-RAG approach in learning discrete latent variable models suggests promising directions for future research, particularly in learning multi-step agents. Basically, RAG can be viewed as a two-step (retrieve-then-generate) agent. Currently, the multi-step trajectory of thinking, reasoning, tool use, and planning in building agents needs to be synthesized or annotated. The JSA-RAG methodology investigated in this paper can be extended to learning such multi-step agents. This avenue of exploration is highly promising and warrants further investigation.
Limitations
Index rebuilding incurs significant computational cost for large KBs. Rebuilding the index scales with the KB size, drastically increasing the training time. In such cases, we need to adopt asynchronous index rebuilding. We leave asynchronous index rebuilding to work with much larger KBs for future work.
@@ Appendix A Training Details
Our overall experiments were run for 20,000 steps and the best results were recorded. During training, we set different learning rates for the retriever and generator. The loss was optimized using the AdamW optimizer. For the dialog task, using GPT-2 and DPR models, the learning rates for the generator and retriever were set to $1\times 10^{-5}$ and $1\times 10^{-5}$, respectively. For the ODQA task with Mistral-7B and BGE models, the learning rates for the generator and retriever were $2\times 10^{-5}$ and $1\times 10^{-5}$, respectively. In the passage concatenation experiment, only the generator was post-trained at a learning rate of $1\times 10^{-5}$ for 10,000 steps. Additionally, the hyperparameter involved in JSA include: MIS sampling steps $m=50$.
@@ Appendix B Computation Cost in Training
To evaluate the computational efficiency of different methods, we measure the training time per 10 iterations on the QuAC dataset (batch size = 1). The results are summarized in Table 8. It can be seen that the training time cost of JSA-RAG is comparable to RAG and VRAG.
@@
Method Time (Seconds / 10 iterations)
JSA-RAG 29 - 30
VRAG 28
RAG 23
Appendix C Gold Passage Annotation
Datasets such as NQ, TQA, and MS-MARCO do not have annotations for gold passages in their corresponding KBs. A gold passage refers to a passage containing information capable of answering a question. To obtain such passages or to find them as closely as possible, we first use the BGE-intialized posterior retriever to retrieve 100 relevant passages from the knowledge base for each question (the posterior retriever performs retrieval using question-answer pairs to incorporate more information). We then prompt GPT-4o to select the one passage that it deems most capable of answering the question from the 100 candidates. The specific prompt is shown in Table 9.
@@ Appendix D LLM Prompt Template
Table 9: Prompt for gold passage selection via GPT-4o
Task Type Prompt Text
Gold Passage Selection
"Question: {question}, Provided Answers: {answers}.
Please select the ID of the passage that best answers
the question from the following paragraphs.
If there is no passage you think can generate the correct answer,
select the ID of the passage that comes closest to answering the question.
Note!!! Only return the value of the passage’s id key."

### v2

@@ 5 Conclusion and Future Work
Remarkably, the potential advantage of the JSA-RAG approach in learning discrete latent variable models suggests promising directions for future research, particularly in learning multi-step agents. Basically, RAG can be viewed as a two-step agent (retrieve-then-generate). Currently, the multi-step trajectory of thinking, reasoning, tool use, and planning in building agents needs to be synthesized or annotated. The JSA-RAG methodology investigated in this paper can be extended to learning such multi-step agents. This avenue of exploration is highly promising and warrants further investigation.
@@ Appendix A Training Details
In our experiments, all methods were run for 20,000 steps with batch size being 1 and the best results were recorded. During training, we set different learning rates for the retriever and generator. The loss was optimized using the AdamW optimizer. For the dialog task, using GPT-2 and DPR models, the learning rates for the generator and retriever were set to $1\times 10^{-5}$ and $1\times 10^{-5}$, respectively. For the ODQA task with Mistral-7B and BGE models, the learning rates for the generator and retriever were $2\times 10^{-5}$ and $1\times 10^{-5}$, respectively. In the passage concatenation experiment, only the generator was post-trained at a learning rate of $1\times 10^{-5}$ for 10,000 steps. Additionally, the hyperparameter involved in JSA include: MIS sampling steps $m=50$.
@@ Appendix B Computation Cost in Training
To evaluate the computational efficiency of different methods, we measure the training time for 100 iteration steps on the OR-QuAC dataset (batch size = 1). The results are summarized in Table 7. It can be seen that the training time cost of JSA-RAG is comparable to RAG and VRAG.
@@
Method Time (Seconds / 100 steps)
JSA-RAG 198
VRAG 153
RAG 148
Appendix C Gold Passage Annotation for retriever evaluation
In testing retrieval performance, some datasets such as NQ, TQA, and MS-MARCO do not have annotations for gold passages in their corresponding KBs. A gold passage refers to a passage containing information capable of answering a question. To obtain such passages or to find them as closely as possible, we first perform a posterior retrieval based on BGE-large to retrieve 100 relevant passages from the knowledge base for each question (the posterior retriever performs retrieval using question-answer pairs to incorporate more information). We then prompt GPT-4o to select the one passage that it deems the most capable of answering the question from the 100 candidates. The specific prompt is shown in Table 8. It should be noted that these GPT-4o-generated annotations are never used in model training, but only used in testing to evaluate the prior retrievers (Table 2) and the posterior retrievers (Table 3), for those datasets without gold passage annotations (i.e., NQ, TQA and MS-MARCO in our experiments).
Table 8: Prompt for Gold Passage Selection via GPT-4o for retriever evaluation
Task Type Prompt Text
Gold Passage Selection Question: {question}, Provided Answers: {answers}. Please select the ID of the passage that best answers the question from the following paragraphs. If there is no passage you think can generate the correct answer, select the ID of the passage that comes closest to answering the question. Note!!! Only return the value of the passage’s id key.